OpenStudy (anonymous):

what are the two solutions to 3x^2-7x-1=0?

OpenStudy (amistre64):

gonna be sqrtish methinks

OpenStudy (anonymous):

x=(7+-sqrt(61))/6

OpenStudy (anonymous):

oh yes I do too.

OpenStudy (amistre64):

(x-7/6)^2 = (1/3)+(49/36) x = \(\frac{7}{6}\pm \frac{\sqrt{61}}{6}\) yep, those fractions always slow me down